OpenStudy (anonymous):

4 3 1 --- x - 3 ---- x = 2 --- 5 4 2 A. -1 9 --- 50 B. 50 - ----- 59 C. 34 ---- 75 D. 9 1 --- 50

OpenStudy (anonymous):

Subtract \[3\frac{ 3 }{ 4 }\] from \[\frac{ 4 }{ 5 }\]

OpenStudy (anonymous):

what do u get?

OpenStudy (anonymous):

50 over 59?

OpenStudy (anonymous):

first turn \[3\frac{ 3 }{ 4}\] into an improper fraction \[\frac{ 15 }{ 4 }\]

OpenStudy (anonymous):

so it becomes\[\frac{ 4 }{ 5 }\] -\[\frac{ 15 }{ 4 }\]

OpenStudy (babyslapmafro):

Note* None of the answers she provided are equal to the variable she is trying to find.

OpenStudy (babyslapmafro):

3(3/4) does not equal 15/4

OpenStudy (babyslapmafro):

3(3/4) = 9/4

OpenStudy (anonymous):

so what one is closer?

OpenStudy (babyslapmafro):

"Closer" is not an option...you need to find an exact answer

OpenStudy (anonymous):

ill just guess lol

OpenStudy (anonymous):

Answ is b, but for the future u need to know how to get there

OpenStudy (babyslapmafro):

If you guess from the choices (a through d) that you have provided, there is a 100% chance that you will be wrong.

OpenStudy (babyslapmafro):

The answer is not B. The answer is -50/29

OpenStudy (anonymous):

can you detail your working to prove it?

OpenStudy (babyslapmafro):



OpenStudy (babyslapmafro):